OpenStudy (anonymous):

pls help! 98. A ball is thrown straight down from the top of a 220-foot building with an initial velocity of -22 feet per second. What is its velocity after 3 seconds? What is its velocity after falling108 feet? position function=-16t^2+v0t+s0 i think the answer to the first part is -118 ft/second and im lost on the second part

OpenStudy (amistre64):

take the derivative to determine velocity

OpenStudy (anonymous):

yea so you would get v(t)=-32t-22 and then you plug in 3 and get -118 correct? what about the second part

OpenStudy (anonymous):

i edited the question sorry it was messed up

OpenStudy (amistre64):

s(t) = -16t^2-22t+220 s'(t) = v(t) = -32t-22, yep. velocity gives speed and direction which tells us that its -118 downward

OpenStudy (amistre64):

solve for t when s(t) = 220-108

OpenStudy (amistre64):

then we can use velocity againt

OpenStudy (anonymous):

so i used the quadratic equation and i got 2, and -27/8

OpenStudy (anonymous):

so now what

OpenStudy (amistre64):

well, since time tends to start at 0 and just grows from there, id let t=2

OpenStudy (amistre64):

v(2) = -32(2) - 22